undeadknight26 (undeadknight26):

The expression 900(1.003274)4x shows the quarterly interest rate Joey earns after x years. Which expression shows the x isolated in the exponent in an equivalent form? 900(1.01316)x 900(4.013096)x 900(0.250819)x 900(1.000817)x

undeadknight26 (undeadknight26):

@skullpatrol @Smexi_Girl

undeadknight26 (undeadknight26):

@april115 @Gatorgirl please help

undeadknight26 (undeadknight26):

@SolomonZelman @RadEn

OpenStudy (solomonzelman):

A is the money earned altogether. \[A= 900(1.003274)^{4x}\]this right? Quaterly compound interest.

undeadknight26 (undeadknight26):

right.

OpenStudy (solomonzelman):

\[ A=900(1.003274)^{4x}~~~~~~~~->~~~~~~~~A= 900\color{red} {(1.003274^4)}^x\]

OpenStudy (solomonzelman):

find the red part please.

OpenStudy (solomonzelman):

It's A. \[(1.003274^4)=900(1.01316)^x\]

undeadknight26 (undeadknight26):

oh lol i got: 1.016477

OpenStudy (solomonzelman):

IDk, I used google calculator.

OpenStudy (solomonzelman):

I am 99.99% sure it's A.

undeadknight26 (undeadknight26):

lol ok thanks again :D

OpenStudy (solomonzelman):

Anytime!